# Membership Inference Attacks and Privacy in Topic Modeling

### v1

@@ Appendix E Existing DP Topic Models
Zhu et al. (2016) [34] $\varepsilon$-DP document-level CGS
Zhao et al. (2021) [33] $\varepsilon$-DP word-level CGS
Huang and Chen (2021) [19] $(\varepsilon,\delta)$-DP word-level CGS sub-sampling
Huang et al. (2022) [18] Rényi-DP word-level CGS
Park et al. (2018) [22] $(\varepsilon,\delta)$-DP document-level VI moments accountant and sub-sampling
Decarolis et al. (2020) [9] Rényi-DP document-level SA local sensitivity from propose-test-release
Wang et al. (2022) [31] Rényi-DP author-level PH pain-free smooth sensitivity
@@
Proof (sketch). First, let $M_{1}$ be a $(\varepsilon_{1},\delta_{1})$-DP vocabulary selection algorithm that returns a private vocabulary set $S$, and $M_{2}$ be a $(\varepsilon_{2},\delta_{2})$-DP topic modeling algorithm that returns a private topic-word distribution $\Phi$. Now, let PRE be a function that takes a corpus $D$ and applies standard pre-processing procedures, and SAN be a function that takes a corpus $D$ and a vocabulary set $S$ and removes all words $w\in D$ if $w\notin S$. Finally, let $M_{x}$ represent the FDPTM algorithm such that for a corpus $D$
@@
Definition F.1 ($k$-Stability [28])
@@
Lemma F.1 (Composition with Stable Functions [28])
@@
Via our definition of $M_{1}$, and using the fact that PRE is a stable function, then the first term in $M_{x}$, $M_{1}(\texttt{PRE}(D))$, satisfies $(\varepsilon_{1},\delta_{1})$-DP. The second term of $M_{x}$ applies $M_{2}$ which directly depends on $M_{1}$ and the stable functions PRE and SAN. Therefore, via adaptive composition, $M_{x}$ is $(\varepsilon_{1}+\varepsilon_{2},\delta_{1}+\delta_{2})$-DP [10]. $\square$

### v2

@@ Appendix E Existing DP Topic Models
Zhu et al. (2016) Zhu et al. (2016) $\varepsilon$-DP document-level CGS
Zhao et al. (2021) Zhao et al. (2021) $\varepsilon$-DP word-level CGS
Huang and Chen (2021) Huang & Chen (2021) $(\varepsilon,\delta)$-DP word-level CGS sub-sampling
Huang et al. (2022) Huang et al. (2022) Rényi-DP word-level CGS
Park et al. (2018) Park et al. (2016) $(\varepsilon,\delta)$-DP document-level VI moments accountant and sub-sampling
Decarolis et al. (2020) Decarolis et al. (2020) Rényi-DP document-level SA local sensitivity from propose-test-release
Wang et al. (2022) Wang et al. (2022) Rényi-DP author-level PH pain-free smooth sensitivity
@@
First, let $M_{1}$ be a $(\varepsilon_{1},\delta_{1})$-DP vocabulary selection algorithm that returns a private vocabulary set $S$, and $M_{2}$ be a $(\varepsilon_{2},\delta_{2})$-DP topic modeling algorithm that returns a private topic-word distribution $\Phi$. Now, let PRE be a function that takes a corpus $D$ and applies standard pre-processing procedures, and SAN be a function that takes a corpus $D$ and a vocabulary set $S$ and removes all words $w\in D$ if $w\notin S$. Finally, let $M_{x}$ represent the FDPTM algorithm such that for a corpus $D$
@@
Definition F.1 ($k$-Stability (Thakurta & Smith, 2013))
@@
Lemma F.1 (Composition with Stable Functions (Thakurta & Smith, 2013))
@@
Via our definition of $M_{1}$, and using the fact that PRE is a stable function, then the first term in $M_{x}$, $M_{1}(\texttt{PRE}(D))$, satisfies $(\varepsilon_{1},\delta_{1})$-DP. The second term of $M_{x}$ applies $M_{2}$ which directly depends on $M_{1}$ and the stable functions PRE and SAN. Therefore, via adaptive composition, $M_{x}$ is $(\varepsilon_{1}+\varepsilon_{2},\delta_{1}+\delta_{2})$-DP (Dwork et al., 2006a). $\square$